OpenStudy (anonymous):

If f(x) = 2x2 − 10, find f(5). i got to 15=2x2 but i cant figure out what to do after

jimthompson5910 (jim_thompson5910):

f(x) = 2x^2 − 10 f(5) = 2(5)^2 − 10 Did you get this far?

OpenStudy (anonymous):

yes then i added 10 to each side

jimthompson5910 (jim_thompson5910):

why? you only need to do that if you're solving for x

jimthompson5910 (jim_thompson5910):

simply evaluate 2(5)^2 − 10

OpenStudy (anonymous):

oh oops

jimthompson5910 (jim_thompson5910):

that's ok

jimthompson5910 (jim_thompson5910):

i know, she wanted to add 10 to both sides, but you only do that when solving for x

OpenStudy (anonymous):

how am i going to find x?

jimthompson5910 (jim_thompson5910):

f(5) means find f(x) when x = 5

jimthompson5910 (jim_thompson5910):

you already know what x is in this case

OpenStudy (anonymous):

so its 5

jimthompson5910 (jim_thompson5910):

f(x) = 2x^2 - 10 so replace EVERY x with 5 to get f(5) = 2(5)^2 - 10

OpenStudy (anonymous):

then 2 times 5?

jimthompson5910 (jim_thompson5910):

So your job now is to find 2(5)^2 - 10 Hint: (5)^2 means 5*5 and 2(5)^2 is 2 times 5 squared

jimthompson5910 (jim_thompson5910):

close, but no

jimthompson5910 (jim_thompson5910):

PEMDAS says you have do exponents before multiplication

OpenStudy (anonymous):

25 then 10?

jimthompson5910 (jim_thompson5910):

f(5) = 2(5)^2 - 10 f(5) = 2(25) - 10 Note: 5^2 = 5*5 = 25

jimthompson5910 (jim_thompson5910):

now multiply 2 and 25 because 2(25) means 2 times 25

OpenStudy (anonymous):

so 50

jimthompson5910 (jim_thompson5910):

then subtract 10

OpenStudy (anonymous):

40

jimthompson5910 (jim_thompson5910):

perfect

jimthompson5910 (jim_thompson5910):

So here's the complete step by step solution f(x) = 2x^2 - 10 f(5) = 2(5)^2 - 10 f(5) = 2(25) - 10 f(5) = 50 - 10 f(5) = 40

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

you're welcome